OpenStudy (anonymous):

What are the zeros of the polynomial function f(x) = x3 – 4x2 – 12x?

OpenStudy (anonymous):

to find zeros of a polynomial we want to factor to get values of x that make the whole thing 0. \[f(x) = x^3 -4x^2 -12x \] what term is in each part?

OpenStudy (anonymous):

a Polynomial of degree 'n' will have n zeros, this is a polynomial of degree 3, therefore there will be 3 values of x that make the whole equation 0.

OpenStudy (anonymous):

0, –2, 6 –2, 6 0, 2, –6 2, –6 here is the answers. & I'm still confused.

OpenStudy (anonymous):

I understand the (X) is the term in each part.

OpenStudy (anonymous):

ok so 'pull out' the 'x'

OpenStudy (anonymous):

\[f(x) = x^3 -4x^2-12x = x(x^2 -4x -12)\]

OpenStudy (anonymous):

which all equals 0

OpenStudy (anonymous):

to find zeros you set f(x) = 0

OpenStudy (anonymous):

So its a?

OpenStudy (anonymous):

lets keep going until you know for sure.

OpenStudy (anonymous):

Yea, cause I am really lost. After you pulled the "x" out i got lost.

OpenStudy (anonymous):

do you get how we can pull out the x?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

we are reversing the distribution process. \[x(a+b) = ax+bx\]

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

here we are starting with the right hand side and going 'backwards'

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

what value of x makes the following true: x*(anything) = 0

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[f(x) = 0 = x(x^2-4x-12)\]

OpenStudy (anonymous):

so in this problem either 'x' = 0 or \[x^2 - 4x -12 = 0\]

OpenStudy (anonymous):

Okay so then how do you get one of these answers? 0, –2, 6 –2, 6 0, 2, –6 2, –6

OpenStudy (anonymous):

to find zeroes of a polynomial what we do is take the polynomial and factor it into seperate pieces of x such as " x(x+a)(x+b) = 0 and then to find what values of x makes this true either of the followings need to occur, if x=0 , if (x+a)=0, or if (x+b) = 0

OpenStudy (anonymous):

the number of zeroes in a polynomial is equal to the number of the power on the highest exponent, in this case 3

OpenStudy (anonymous):

so we are looking for 3 components that any part can equal 0 to make the whole function = 0

OpenStudy (anonymous):

So b and d are out.

OpenStudy (anonymous):

yes, for this one.

OpenStudy (anonymous):

the terms of x(x+a)(x+b) are called 'roots'

OpenStudy (anonymous):

Okay, then what do we do.

OpenStudy (anonymous):

they correspond to where the function, f(x), crosses the x axis

OpenStudy (anonymous):

so we have 1 root already, that is x=0

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[x(x^2-4x-12) = 0\] this is when either x = 0 or when \[x^2-4x-12 = 0\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

since the order of this is 2 we need two values of x that make this true

OpenStudy (anonymous):

So its a.

OpenStudy (anonymous):

\[x^2-4x-12 = (x+a)(x+b)\]

OpenStudy (anonymous):

-12 = a*b -4x = a*x+b*x x^2 = x*x

OpenStudy (anonymous):

since x is in all terms of the 2nd equation you can divide the whole thing by 'x' and not cahnge it so it becomes: -4 = a + b

OpenStudy (anonymous):

so you have: -12 = a*b and -4 = a+b

OpenStudy (anonymous):

what will a and b equal?

OpenStudy (anonymous):

-2 and 6 right?

OpenStudy (anonymous):

-4 = -2 + 6?

OpenStudy (anonymous):

now plug a and b into the equation above to get \[x^2 -4x -12 = (x-6)(x+2) = 0\]

OpenStudy (anonymous):

now what values of 'x' makes each root = 0?

OpenStudy (anonymous):

x = 0 or -2 or 6 to make f(x) = 0